Spain Sudan Plastic Opp Cost of Plastic Gold Opp Cost of gold Plastic Opp cost of...

Spain

Sudan

Plastic

Opp Cost of Plastic

Gold

Opp Cost of gold

Plastic

Opp cost of Plastic

Gold

Opp cost of gold

A

0

---------

160

0

---------

280

B

360

120

420

210

C

720

80

840

140

D

1080

40

1260

70

E

1440

0

---------

1680

0

--------

Based on the table above

A

Sudan has the comparative advantage in plastic

B

Sudan has the comparative advantage in gold

C

Sudan should specialize in plastic

D

Sudan has tariffs on plastics
Answer #1

Spain's opportunity cost of plastic = 160/1440 = 0.11 of Gold.

Spain's opportunity cost of Gold = 1440/160 =9 of Plastic.

Sudan's opportunity cost of plastic = 280/1680 = 0.167 of Gold

Sudan's opportunity cost of Gold = 1680/280 = 6 of plastic.

Because Sudan has lower opportunity cost for the production of plastic. This implies that Sudan has the comparative advantage in plastic. Hence,option(A) is correct.